OpenStudy (anonymous):

the triangle on the left has a perimeter of 14. while on the right the triangle has a perimeter of 21. what are x and y 2y+x=14 2(5/4 y) +3x=21

OpenStudy (anonymous):

@winterfez

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

actualy y=-6 x=26

OpenStudy (anonymous):

if you plug in \[2(-6)+26=14\]

OpenStudy (anonymous):

but how'd you get those answers because i solved for y in the first equation and then plugged it in to the second one.

OpenStudy (anonymous):

wait sorry x not 26

OpenStudy (anonymous):

ya you right y=6 x=2

OpenStudy (anonymous):

:)))) thanks

OpenStudy (anonymous):

is \[\frac{ 5 }{ 4y }\] or \[\frac{ 5 }{ 4 }y\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

oh ok..your answer is right

OpenStudy (anonymous):

i just dont know how to use the equation part to get a regular fraction

OpenStudy (anonymous):

thanks so much :)

OpenStudy (anonymous):

np..try to get rid of decmical point and it will be easier to find the fraction

OpenStudy (anonymous):

ok :)